OpenStudy (anonymous):

Medal for help :) 2/3x = 6 A. x= 9 B. x= 4 C. x=18 D. x= 12

OpenStudy (anonymous):

my answer is a. x=9

jimthompson5910 (jim_thompson5910):

(2/3)*x = 6 (2/3)*9 = 6 6 = 6 So that checks out